A psychology professor assigns letter grades on a test according to the following scheme. A: Top 8% of scores B: Scores below the top 8% and above the bottom 60% C: Scores below the top 40% and above the bottom 18% D: Scores below the top 82% and above the bottom 9% F: Bottom 9% of scores Scores on the test are normally distributed with a mean of 67.7 and a standard deviation of 8.5. Find the numerical limits for a D grade. Round your answers to the nearest whole number, if necessary.
Answer #1

z score for bottom 18% area = -0.92

z score for bottom 9% area = -1.34

Hence,

Lower limit for D grade = 67.7 - 1.34*8.5 = 56

Upper limit for D grade = 67.7 - 0.92*8.5 = 60